OpenStudy (anonymous):

Finding the x int of: (x^2-1)/(x^2-2x-3)

OpenStudy (anonymous):

I know that you set it equal to zero and solve. But I'm having trouble solving it.

OpenStudy (amistre64):

when does a fraction equal zero?

OpenStudy (anonymous):

So your saying there is no x-int?

OpenStudy (amistre64):

no, im saying ... when does a fraction equal zero?

OpenStudy (anonymous):

when zero is in the numerator?

OpenStudy (amistre64):

correct, what possible values does that give us for this? keeping in mind the the bottom cant zero out either

OpenStudy (amistre64):

when does x^2 - 1 = 0 ?

OpenStudy (anonymous):

would it be only positive 1?

OpenStudy (amistre64):

well, (+1)^2 = 1, and (-1)^2 = 1 check if 1 and -1 dont make the bottom go zero, if not then they are good

OpenStudy (amistre64):

if x^2-2x-3 = 0 when x=1 ; then we cant include 1 as a solution if x^2-2x-3 = 0 when x=-1 ; then we cant include -1 as a solution otherwise, we keep them as solutions

OpenStudy (amistre64):

do you see why? when the bottom of a fraction is zero, its undefined

OpenStudy (anonymous):

i just checked and both -1 and 1 do not make the denominator =0

OpenStudy (anonymous):

so they both will be the x int?

OpenStudy (amistre64):

x^2-2x-3 -1 -1 ------------ 1 + 2- 3 = 0 ; -1 does make a zero underneath

OpenStudy (amistre64):

the x=1 makes a -4 i believe so its the only good answer

OpenStudy (anonymous):

your right...i messed up

OpenStudy (amistre64):

the graph should verify it :) http://www.wolframalpha.com/input/?i=%28x%5E2-1%29%2F%28x%5E2-2x-3%29

OpenStudy (anonymous):

thanks for the help!

OpenStudy (amistre64):

good luck :)